OpenStudy (anonymous):

1. square root of 6 divided by square root of 27? 2.) 4 divided by square root of 5? 3.) square root of 5 divided by square root of 3? 4) square root of 3x squared y to the 3d power. please help. i have so many questions. i dont understand math.

OpenStudy (anonymous):

1. \[\frac{ \sqrt{6} }{ \sqrt{27} } = \sqrt{\frac{ 6 }{ 27 }}\]

OpenStudy (anonymous):

The threes cancel, so we have \[\sqrt{\frac{ 2 }{ 9 }}\] Then we convert it back to the original form, \[\frac{ \sqrt{2} }{ \sqrt{9} }\] sqrt 9 is 3, so our final answer is \[\frac{ \sqrt{2} }{ 3 }\]

OpenStudy (anonymous):

2. I believe that all you have to do here is rationalize the denominator, so we mulitply the top and the bottom by sqrt5/sqrt5. We can do this because it is equivalent to 1. \[\frac{ 4 }{ \sqrt{5} } * \frac{ \sqrt{5} }{ \sqrt{5} } = \frac{ 4\sqrt{5} }{ 5 }\] and we are done.

OpenStudy (anonymous):

3. This is the same as number 2, so we rationalize the denominator and multiply by sqrt3/sqrt3. \[\frac{ \sqrt{5} }{ \sqrt{3} } * \frac{ \sqrt{3} }{ \sqrt{3} } = \frac{ \sqrt{15} }{ {3} }\] and we are done.

OpenStudy (anonymous):

for number 4, could you restate the question in the equation form? I'm not quite sure what you are asking, did you mean \[\sqrt{3 x^2y}^ { 3}\]